The common stock of COB, Inc. is currently selling at $80 per share. The directors wish to reduce the share price and increase the share volume prior to a new issue.


Section G--Stock Split and Stock Dividend 

The common stock of COB, Inc. is currently selling at $80 per share. The directors wish to reduce the share price and increase the share volume prior to a new issue. The per share par value is $10.500,000 shares are issued and outstanding. 

Prepare the necessary journal entries assuming the following. 

 1. The board votes a 2 for 1 stock split. 

2. The board votes a 100% stock dividend. 

3. After the above actions, what is the anticipated selling price of COB common stock.
Answer #1

1)

No Journal entry required for stock split .It requires only book adjustment where number of shares is increased to 500000*2/1 = 1000000 shares and thereby reducing par value to 10*1/2 =$ 5per share

2)

Account title Debit credit
Stock dividend /Retained earning 5000000
Common stock (500000*10) 5000000

Number of shares issued as stock dividend= 500000*100% = 500000

In case of large stock dividend (more than 25% ),Amount of stock dividend is equals to par value of shares issued.

3)

Market price per share
Stock Split Market value per share is reduced by stock split 80 *1/2 =$ 40 per share
Stock dividend Market capitalization /number of shares outstanding after stock dividend 40000000/1000000 =$ 40 per share

Working"

Market capitalization before stock dividend= 500000*80= 40000000

Number of shares outstanding after stock dividend = 500000+500000= 1000000
Answer #2
1 and 2
Date Particulars Debit Credit
XXXX Common stock @10 Per share…………………Dr      5,000,000
    To Common stock @ 5 Per share      5,000,000
( Being 500000 shares are split into 1000000 shares)
XXXX Retained Earnings A/c………………………..Dr    80,000,000
     To Common stock      5,000,000
      To Excess Capital value    75,000,000
( Being 100% stock dividend issued)

Note for second entry

Particulars Amount Explanation
Original number of shares              500,000
number of shares after 2 for one stock split          1,000,000 2 shares for one shares held hence number of shares are doubled
100 % stock dividend therefore new 10 Lakh shares are issued          1,000,000
Par Value of stock dividend          5,000,000 10 Lakhs shares X 5 Par Value
Additional capital        75,000,000 10 Lakhs shares X 75 Par Value (80 - 5)

3) After the above both the actions the total number of shares increased from 500,000 shares to 2,000,000 shares.

The Market selling price theoretically is calculated by Total market value of the company divided by number of shares outstanding. As the number of shares outstanding increase therefore the value of selling price will be reduced in future of the company.